OpenStudy (anonymous):

How do you take the antiderivative of (2x-1)^-1/2???

OpenStudy (anonymous):

Is that (2x-1)^(-1/2) or (2x-1)^-1 which is divided by 2?

OpenStudy (anonymous):

(2x-1)^(-1/2)

OpenStudy (anonymous):

O.k. Then you would let u = 2x-1 Then du = 2 dx So your integral:\[\int\limits \frac{1}{\sqrt{2x+1}}dx\] becomes

OpenStudy (anonymous):

1/2 times \[\int\limits \frac{2dx}{\sqrt{2x+1}}\]

OpenStudy (anonymous):

\[\int\limits \frac{du}{\sqrt{u}}=\]

OpenStudy (anonymous):

\[\int\limits u ^{-1/2}du=\]

OpenStudy (anonymous):

\[\frac{u ^{-1/2 + 1}}{-1/2+1}+C\]

OpenStudy (anonymous):

\[\frac{(2x+1) ^{1/2}}{1/2}+C\]

OpenStudy (anonymous):

and dividing by 1/2 is the same as multiplying by the reciprocal 2/1

OpenStudy (anonymous):

What about x/(2x-1)^(1/2) with an upper of 5 and a lower of 1. My calculator is giving me a correcct answer of 5.333333. But I can work the problem out to be the same

OpenStudy (anonymous):

sorry, I "can't" work the problem out the same

OpenStudy (anonymous):

o.k. well, you would want to plug 5 into \[2*\sqrt{2x+1}\] and then subtract what you get when you plug 1 into it.

OpenStudy (anonymous):

that would be 2 times the result of \[\sqrt{2(5)+1} - \sqrt{2(1)+1} = \sqrt{11}-\sqrt{3}\]

OpenStudy (anonymous):

which turns out to be about 3.17

OpenStudy (anonymous):

Oh! I didn't notice you had an "x" in the numerator.

OpenStudy (anonymous):

You would still use the same substitution

OpenStudy (anonymous):

but the x in the numerator can be rewritten in terms of "u"

OpenStudy (anonymous):

right. x= (1+u)/2

OpenStudy (anonymous):

since u = 2x+1 u-1 = 2x and so you get that x = (.5)*(u - 1)

OpenStudy (anonymous):

so dividing that by \[\sqrt{u} = u^{1/2}\] you can use exponent rules to get

OpenStudy (anonymous):

\[\frac{.5*(u-1)}{u^{1/2}} \]

OpenStudy (anonymous):

which boils down to \[.5* (u^{1-1/2}-u^{-1/2})\]

OpenStudy (anonymous):

which is\[.5* (u^{1/2}-u^{-1/2})\]

OpenStudy (anonymous):

and you can use the power rule for integrals to antidifferentiate

OpenStudy (anonymous):

\[.5* (u^{3/2}-u^{1/2})\]

OpenStudy (anonymous):

or \[.5* ((2x+1)^{3/2}-(2x+1)^{1/2})\]

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

Yes, but when I plug in the upper and lower I'm still not getting the correct answer.......

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

The (2x+1) is suppose to be (2x-1)

OpenStudy (anonymous):

right...it's 14.8511...

OpenStudy (anonymous):

Oh! :) Well, I can guarantee you if you use 2x-1 instead it should work but I'll check

OpenStudy (anonymous):

My graphing calculator is giving me 5.333333

OpenStudy (anonymous):

\[9^{3/2}-9^{1/2}-(1^{3/2}-1^{1/2})\]

OpenStudy (anonymous):

is \[3\sqrt{3}-3\]

OpenStudy (anonymous):

Do you have a graphing cal on you?

OpenStudy (anonymous):

Actually, you have to go back and change the substitution also so it is more work than this

OpenStudy (anonymous):

if u = 2x-1 then x=.5*(u+1)

OpenStudy (anonymous):

Thanks for your help. I'm in FLorida and it's about 1am. I need to hit the hay. Thanks again......

OpenStudy (anonymous):

Yeah, I did it by hand and it works out. Just use that substitution

OpenStudy (anonymous):

and also multiply by the reciprocal of the new exponent you get when you integrate. ;)

OpenStudy (anonymous):

Will do. Thanks!

OpenStudy (anonymous):

Your welcome...sorry I'm new to typing this out. It's a bit distracting ;) but I'll get used to it. Night!

OpenStudy (anonymous):

Yeah, it's a pain in the retrice Laters